### v2

# MedAide: Towards an Omni Medical Aide via Specialized LLM-based Multi-Agent Collaboration

Abstract
Large Language Model (LLM)-driven interactive systems currently show potential promise in healthcare domains. Despite their remarkable capabilities, LLMs typically lack personalized recommendations and diagnosis analysis in sophisticated medical applications, causing hallucinations and performance bottlenecks. To address these challenges, this paper proposes MedAide, an LLM-based omni medical multi-agent collaboration framework for specialized healthcare services. Specifically, MedAide first performs query rewriting through retrieval-augmented generation to accomplish accurate medical intent understanding. Immediately, we devise a contextual encoder to obtain intent prototype embeddings, which are used to recognize fine-grained intents by similarity matching. According to the intent relevance, the activated agents collaborate effectively to provide integrated decision analysis. Extensive experiments are conducted on four medical benchmarks with composite intents. Experimental results from automated metrics and expert doctor evaluations show that MedAide outperforms current LLMs and improves their medical proficiency and strategic reasoning.
MedAide: Towards an Omni Medical Aide via Specialized
LLM-based Multi-Agent Collaboration
[…]
1 Introduction
Developing the goal-oriented conversation system Lee et al. (2024); Cheng et al. (2023) has received increasing attention in recent years. Advanced Large Language Models (LLMs) OpenAI (2022); Achiam et al. (2023); OpenAI (2024) have shown exceptional generalization capabilities in general scenarios for multiple purposes within human-machine interactions, such as negotiation He et al. (2018) and persuasion Wang et al. (2019). In this context, LLM-centered interactive medical assistants Bao et al. (2023); Yang et al. (2024a, 2023b); Shi (2023); Chen et al. (2023b) have become research hotspots that promise to improve diagnosis efficiency and promote service automation. Previous attempts infuse LLMs with healthcare-specific knowledge through relevant corpus construction (e.g., knowledge databases Li et al. (2023b)) and multi-stage training procedures (e.g., supervised fine-tuning Xiong et al. (2023)). Although these strategies stimulate the models’ comprehension of medically relevant intents, there are still bottlenecks when faced with real-world applications that require sophisticated reasoning.
Given the mimicking of learned feedback in human behaviors Du et al. (2023); Park et al. (2023), automated medical agent construction promises to enhance LLMs’ instruction following and logical profiling capabilities Lee et al. (2024). The collaboration among multiple agents to handle different patient inquiries and symptomatic caseloads facilitates accurate dialog goal fulfillment while accounting for individual differences Fan et al. (2024); Li et al. (2024). Despite significant advancements, current efforts focus primarily on medical education training Wei et al. (2024) or selective question-answering Tang et al. (2024), lacking understanding of complicated intents behind user queries and giving hierarchical therapeutic recommendations.
To address these issues, this paper proposes an LLM-based omni medical agent framework for real-world medical applications, called MedAide. The core philosophy of MedAide lies in first complementing and decomposing multi-dimensional medical intents through query rewriting to enhance the model’s dissection of composite medical queries. Then, we introduce a contextual encoder to learn intent prototype embeddings from multifaceted instructions, which is utilized to specify fine-grained target intents via the similarity matching lemma. In this case, different paramedical agents are correspondingly activated to provide personalized determinations with specialized medical expertise. Ultimately, we formulate a decision analysis module with Chain-of-Thought (CoT) properties to summarize the responses of activated agents in a faithful manner to make integrated decisions. The main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to propose the omni multi-agent collaboration framework for real-world scenarios with composite healthcare intents, which shows potential for advancing interactive systems for personalized healthcare.
  • Our MedAide effectively improves the strategic reasoning of LLMs under complicated dialog goals through feedback and collaboration of specialized paramedical agents.
  • Extensive experiments on seven medical benchmarks spanning 17 types of rich intents prove the effectiveness of MedAide. As a plug-and-play framework, MedAide can be readily combined with current LLMs and provides competitive improvements.
2 Related Work
2.1 LLMs in Healthcare Domains
The Large Language Models (LLMs) exemplified by ChatGPT (OpenAI, 2022) and GPT-4 Achiam et al. (2023) demonstrate excellent performance in multidisciplinary applications. Although current LLMs Bai (2023); Yang et al. (2023a); AI (2024) with certain medical knowledge benefit from large-scale corpus support, they lack specialized medical proficiency and have significant performance bottlenecks in domain-specific scenarios. Recently, several attempts Chen et al. (2023b, a); Xu (2023); Bao et al. (2023); Yang et al. (2024a) have begun to build medically customized LLM assistants to fulfill the diagnostic and consultative demands. For instance, HuatuoGPT series Zhang et al. (2023); Chen et al. (2023a) have shown promising results in bridging generic-medical knowledge gaps by absorbing real doctor-patient conversations. ZhongJing series Yang et al. (2023b); Shi (2023) improve Chinese medical capabilities by introducing expert feedback and multi-round medical instructions. In addition, PediatricsGPT Yang et al. (2024a) proposes a systematic training framework to construct interactive healthcare systems for pediatric specialists and medical generalists. Unlike previous studies, our framework aims to more fully recognize medical intents and refine the models’ reasoning abilities in complex scenarios through the LLM-based multi-agent collaboration.
2.2 LLM-based Multi-agent Collaboration
With the focus of researchers on sophisticated goal-oriented dialog generation Wu et al. (2023); Fan et al. (2024); Jiang et al. (2024), the inherent dilemmas of hallucinatory responses and weak comprehension Chen et al. (2024); Yang et al. (2024b) in LLMs have been gradually exposed. In this context, LLM-based automated agents are proposed to provide effective perception and decision-making skilfulness by incorporating external tools and databases Cai et al. (2023); Li et al. (2023a). By mimicking human behavioral logic, multiple agents perform feedback and collaboration to enhance diverse intent understanding tasks, including educational training Lee et al. (2024); Wei et al. (2024), emotional comfort Cheng et al. (2023), and workflow integration Hong et al. (2023). For example, MEDCO Wei et al. (2024) enables LLMs to simulate patients and doctors to enhance the practice performance of virtual students in interactive environments. MedAgents Tang et al. (2024) improves the performance of medical assistants in zero-shot settings through the role-playing strategy. In comparison, the proposed MedAide focuses more on mining profound healthcare intents and moving towards robust healthcare practices.
3 Methodology
Figure 1 presents the overall procedure of the proposed MedAide. Here, we detail the systematic workflow procedure, including query rewriting, intent recognition, and agent collaboration.
Figure 1: Illustration of the proposed MedAide framework. MedAide consists of three major phases: (i) query rewriting aims to reformulate medical queries and mine user intents to ensure the system accurately understands the input requirement goals. Among the optimized queries, (ii) intent recognition utilizes a designed contextual encoder to semantically match distinct medical intents to achieve accurate intent recognition. Based on the recognition results, (iii) agent collaboration dynamically activates the corresponding paramedical agents. Then, a decision analysis module incorporates the outputs of multiple agents and generates integrated analysis results.
3.1 Query Rewriting
In this phase, we first process the initial query through the query input processor. The processor is based on a set of syntactic regularization algorithms that combine LLMs with a predefined ruleset $R$ (please refer to Appendix A.1 for details). The core concept is to check, optimize, and normalize user input to $Q_{std}$. Algorithm 1 shows the corresponding program. Then, a key element extractor distills critical information from $Q_{std}$, such as symptoms, condition descriptions, and medical histories, to form the element set $E_{i}$. In the Retrieval-Augmented Generation (RAG), we build an indexed database of 1,095 expert-reviewed medical guidelines and retrieve documents related to these elements using a semantic retrieval method  Lewis et al. (2021), forming the document set $D_{ref}$. After that, the recall analysis module selects the conforming documents by prompting guidance, which is fed into the LLM-based query prompter along with $Q_{std}$. The prompter optimizes the empirical information and efficient decomposition of composite intents. Eventually, we design a refined query constructor to merge and integrate the generated subqueries. It filters and reconstructs multiple subqueries $Q_{gen}$ with a predefined ruleset (please refer to Appendix A.2 for details) to ensure semantically complete and formally uniform output.
Algorithm 1 LLM-Based Input Standardization
  Input: User query $Q_{input}$, Large Language Model $\mathcal{M}$, Set of rules $\mathcal{R}$
  Output: Standardized query $Q_{std}$
  Initialize: $Q_{cur}\leftarrow Q_{input}$,$converged\leftarrow\text{False}$
  while not $converged$ do
     $converged\leftarrow\text{True}$
     for each rule $r\in\mathcal{R}$ do
        $Q_{new}\leftarrow\mathcal{M}(Q_{cur},r)$
        if $Q_{new}\neq Q_{cur}$ then
           $Q_{cur}\leftarrow Q_{new}$
           $converged\leftarrow\text{False}$
        end if
     end for
  end while
  return $Q_{std}\leftarrow Q_{cur}$
3.2 Intent Recognition
After the query rewriting, the optimized query $Q_{opt}$ is matched with a set of intent prototype embeddings $E_{i}$ generated by a contextual encoder, which is designed to capture the semantic features of different medical intents. The encoder is constructed on top of BioBERT Lee et al. (2019) to learn prototype representations by performing fine-grained intent classification. We add a fully connected layer after the embedding layer with output dimensions aligned to the 17 medical intent categories and generate the corresponding probability distributions via a softmax activation function. Specifically, the contextual encoder maps the optimized query together with the intent embeddings into a 768-dimensional embedding space. It computes the cosine similarity $S_{ij}$ between the query and each intent embedding $E_{i}$ with the following formula:
$S_{ij}=\frac{Q_{opt}\cdot E_{i}}{\left\|Q_{opt}\right\|\left\|E_{i}\right\|}.$ (1)
Subsequently, the probability distribution $\alpha_{ij}$ for each intent after the softmax is expressed as:
$\alpha_{ij}=\frac{\exp(S_{ij})}{\sum_{l=1}^{17}\exp(S_{il})}.$ (2)
If the probability $\alpha_{ij}$ of an intent $i$ exceeds a predetermined threshold, the intent will be activated, triggering the corresponding agent:
$\text{Activated Intent}_{i}=\left\{\begin{array}[]{ll}1&\text{if }\alpha_{ij}>\text{Threshold},\\ 0&\text{otherwise}.\end{array}\right.$ (3)
In this way, our framework can automatically activate the most compliant medical intent based on the optimized query, directing it to the corresponding agent to perform subsequent operations.
3.3 Agent Collaboration
The corresponding agents are dynamically activated when the framework recognizes relevant and valid intents. Considering comprehensiveness and diversity, agent collaboration covers systematic healthcare services spanning pre-diagnosis, diagnosis, medicament, and post-diagnosis. Also, an LLM-based decision analysis module is employed for information integration and summarization.
Specifically speaking, the pre-diagnosis agent equips the patient repository, which is a relational database Codd (1970) for storing historical patient visit records. These records can be utilized for health assessments, identifying potential risks, and recommending appropriate departments to visit.
The diagnosis agent incorporates 506 high-quality cases provided in the medical record database Fan et al. (2024). We employ a hybrid retrieval scheme Sawarkar et al. (2024) to obtain demonstrations similar to the patient’s condition, thus providing reliable evidence regarding precise treatment recommendations in symptom analysis and etiological ranking intents. The hybrid retrieval consists of two main components: keyword retrieval and vector retrieval. First, the framework extracts a subset of documents $D_{\text{slice}}$ containing the query keyword from the document collection $D$ by the keyword retrieval of Eq. (1):
$D_{\text{slice}}=\{d\in D\mid\text{KeywordMatch}(Q,d)=True\}.$ (4)
Afterward, MedAide calculates the similarity $S(Q,d)$ between the query $Q$ and the document $d$ according to Eq. (2) by vector retrieval and retains the documents whose similarity is higher than a threshold set $\tau$, forming a subset $D_{\text{match}}$:
$S(Q,d)=\frac{E_{Q}\cdot E_{D}}{\|E_{Q}\|\|E_{D}\|},$ (5)
$D_{\text{match}}=\{d\in D\mid S(Q,d)>\tau\}.$ (6)
Subsequently, the keyword matching results in $D_{\text{slice}}$, and the vector retrieval results in $D_{\text{match}}$ are merged to obtain the final set of relevant documents $D_{\text{final}}$. These documents will provide empirical support for diagnostic recommendations:
$D_{\text{final}}=D_{\text{slice}}\cup D_{\text{match}}$ (7)
In this case, the diagnosis agent further interprets the ancillary examination results to help patients understand their health status. Meanwhile, it provides a second diagnostic opinion to minimize the risk of misdiagnosis.
The medicament agent utilizes 26,684 medication data extracted from the PubMed library Pud (2024) and provides patients with comprehensive medication counseling through keyword and semantic hybrid retrieval. The holistic counseling includes drug usage, dosage recommendations, potential drug interactions, and contraindications assessments to ensure the safety and efficacy of the medication. In addition, it helps patients understand potential side effects and suggests the best timing for medication administration to optimize treatment outcomes.
The post-diagnosis agent focuses on providing personalized rehabilitation advice and emotional support. It tracks and manages the recovery process by updating the patient repository with patient information.
The Decision Analysis Module (DAM) with the CoT properties combines the output of specific activated agents with the patient’s medical histories to dynamically adjust the treatment protocols and formulate integrated rehabilitation plans. The final answer generated by DAM based on $D_{\text{final}}$, the prompt, and the agent output $A_{i}$ is expressed as follows:
$Answer=\text{DAM}(A_{i},\text{prompt},D_{\text{final}}).$ (8)
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 9.06/9.18 2.68/2.81 17.54/19.83 1.70/2.86 10.04/11.74 3.66/4.19
Meditron-7B/ + MedAide 3.76/4.94 0.91/1.30 9.63/10.51 0.64/0.68 4.52/5.21 1.74/2.07
HuatuoGPT-II/ + MedAide 14.6/15.18 5.87/6.63 26.81/31.84 5.26/9.88 13.57/19.04 6.43/7.25
Baichuan4/ + MedAide 12.97/15.95 5.39/7.85 22.68/33.60 4.68/11.69 13.33/20.75 5.57/7.98
LLama-3.1-8B/ + MedAide 10.95/15.36 2.13/6.78 22.47/32.54 2.62/9.41 11.45/19.29 4.37/6.45
GPT-4o/ + MedAide 15.28/15.93 6.33/7.56 26.89/30.78 5.95/8.65 14.15/17.64 6.34/7.31
Table 1: Comparison results on the Pre-Diagnosis benchmark.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 10.50/11.72 4.80/5.58 24.08/24.59 4.93/4.79 12.94/11.96 5.85/5.79
Meditron-7B/ + MedAide 12.01/13.93 3.13/6.03 19.10/26.53 3.60/6.12 8.60/11.67 3.99/6.48
HuatuoGPT-II/ + MedAide 22.65/25.26 10.21/12.11 34.51/40.32 8.36/11.73 15.77/19.11 8.95/10.65
Baichuan4/ + MedAide 22.58/25.51 10.98/13.61 36.26/42.77 10.30/15.00 17.03/22.43 9.61/11.95
LLama-3.1-8B/ + MedAide 16.92/22.75 6.56/9.97 26.39/38.52 4.78/10.41 12.79/18.68 6.21/8.94
GPT-4o/ + MedAide 21.67/26.80 9.73/14.56 32.67/44.11 7.95/15.81 16.20/23.67 8.41/12.69
Table 2: Comparison results on the Diagnosis benchmark.
4 Experiments
4.1 Datasets and Implementation Details
All experiments are conducted in the zero-shot setting. To reflect the real-world demands for healthcare services, we conduct four different benchmarks across 17 types of medical intents, including Pre-Diagnosis, Diagnosis, Medicament, and Post-Diagnosis benchmarks. Each benchmark contains 500 composite intent instances. Please refer to Figures 7,  8,  9, and  10. in the Appendix for the specific intent taxonomy. For a comprehensive overview of the prompts used during the query rewriting and agent collaboration phases within MedAide, refer to Appendix  B. Additionally, to enable fine-grained medical intent recognition, we develop a contextual encoder training dataset consisting of 2,800 training samples, 300 validation samples, and 300 testing samples. Further configuration details can be found in Appendix A.3.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 11.83/11.91 4.48/4.57 21.61/24.23 4.65/5.91 13.55/14.75 5.32/6.38
Meditron-7B/ + MedAide 5.76/19.43 2.04/17.50 12.29/34.84 2.12/29.43 6.31/32.57 2.79/24.47
HuatuoGPT-II/ + MedAide 17.14/37.89 9.61/27.47 30.72/58.06 11.96/33.62 19.96/45.39 9.57/24.23
Baichuan4/ + MedAide 14.64/44.63 7.12/36.34 26.26/62.11 8.09/42.56 15.47/49.91 7.08/35.99
LLama-3.1-8B/ + MedAide 13.80/23.87 5.90/16.75 27.28/47.39 7.50/26.75 16.05/38.30 6.37/15.77
GPT-4o/ + MedAide 16.23/45.13 8.16/38.96 29.24/65.09 9.78/40.53 17.53/52.71 8.61/37.09
Table 3: Comparison results on the Medicament benchmark.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 12.97/19.91 4.86/6.86 25.31/31.98 4.21/4.48 12.44/15.00 5.63/7.22
Meditron-7B/ + MedAide 8.72/12.78 2.46/3.06 18.75/20.21 1.77/4.32 7.28/8.58 3.84/4.40
HuatuoGPT-II/ + MedAide 21.39/26.10 12.24/12.46 35.55/42.23 11.02/13.62 13.62/21.18 10.59/10.96
Baichuan4/ + MedAide 15.89/26.10 7.06/13.92 27.69/44.55 6.92/13.05 14.86/21.35 6.59/11.61
LLama-3.1-8B/ + MedAide 21.85/22.95 7.42/7.68 35.36/39.71 5.22/7.97 15.02/17.93 7.65/8.01
GPT-4o/ + MedAide 19.27/26.83 9.28/11.54 30.72/40.83 8.64/8.71 16.22/18.12 8.41/9.89
Table 4: Comparison results on the Post-Diagnosis benchmark.
Figure 2: Response comparisons of MedAide with other baselines via doctor evaluation.
4.2 Model Zoo
We compare a series of state-of-the-art (SOTA) models for comprehensive evaluations. Among medical LLMs, ZhongJing2 Shi (2023) is a traditional Chinese medicine model based on Qwen-1.5-1.8B with complete training procedures. Meditron-7B Chen et al. (2023b) performs medically relevant continuous pre-training on Llama-2-7B to extend the breadth and depth of the model in medical knowledge. HuatuoGPT-II (13B) Chen et al. (2023a) employs a one-stage unified approach for domain adaptation to improve medical expertise and model applicability. In general-purpose models, Baichuan4 AI (2024) optimizes long texts by scaling law and combines reinforcement learning techniques to improve reasoning ability. Llama-3.1-8B Dubey et al. (2024) relies on the grouped query attention mechanism to enhance the inference efficiency and performs well on multilingual tasks. GPT-4o OpenAI (2024) shows excellent language comprehension and generation capabilities and excels in handling complex tasks.
4.3 Comparision with SOTA Methods
As a plug-and-play framework, we combine MedAide with the baseline models to provide comprehensive evaluations by different metrics, including BLEU-1/2 (%), ROUGE-1/2/L (%), and GLEU (%). Comparison results on the four medical benchmarks are shown below.
Results on Pre-Diagnosis Benchmark. As Table 1 shows, our framework consistently improves the performance across metrics for all models. (i) For instance, MedAide-based HuatuoGPT-II and Baichuan4 show significant relative improvements in ROUGE-L scores of 40.3% and 55.6%, respectively, suggesting multidimensional health risk assessments for patient cases. (ii) In addition, MedAide helps general LLama-3.1-8B to enhance the diversity and precision of diagnostic content generation, as reflected in the average 79.15% increase of BLEU-1/2 relative to the vanilla baseline. (iii) Despite the performance constraints of ZhongJing2 and Meditron-7B in tackling complex medical scenarios due to the scaling law, our framework still provides sustained performance gains.
Results on Diagnosis Benchmark. Table 2 shows the diagnosis task results containing six composite intent scenarios. (i) Combined with our framework, Meditron-7B and HuatuoGPT-II outperform vanilla baselines in response accuracy and completeness, as evidenced by significant improvements across metrics. These gains show that MedAide assists medical LLMs in giving more comprehensive symptom analyses and treatment suggestions with enhanced hierarchical interpretation capabilities regarding examination results. (ii) Meanwhile, the general models (e.g., LLama-3.1-8B and GPT-4o) equipped with MedAide enhance the specialization and adaptability in healthcare services. (iii) We observe slight performance drops (less than 1%) in ZhongJing-2 on ROUGE-2/L and GLEU. The potential reason is that the model has difficulty in the query rewriting phase to grasp the multifaceted knowledge of the retrieval-augmented generation due to the limited capacity, causing sub-optimal prevention guidelines and second opinions.
Pre-Diagnosis Diagnosis Medicament Post-Diagnosis
Components ROUGE-L GLEU ROUGE-L GLEU ROUGE-L GLEU ROUGE-L GLEU
Full MedAide 17.64 7.31 23.67 12.69 52.71 37.09 18.12 9.89
w/o Query Rewriting 16.02 6.44 18.35 10.37 32.06 32.45 16.48 8.79
w/ GPT-4o Recognition 16.58 6.73 22.98 12.34 44.72 34.10 17.56 9.02
w/o Decision Analysis 17.22 7.22 23.14 12.54 48.72 35.25 18.04 9.56
Table 5: Ablation study results on four benchmarks.“w/” and “w/o” are short for with and without, respectively.
Benchmarks MedAide w/o QR w/ GPT-4 R
Pre-Diagnosis 0.76 0.62 0.47
Diagnosis 0.83 0.61 0.49
Medicament 0.80 0.49 0.51
Post-Diagnosis 0.63 0.56 0.37
Intent Aggregation 0.86 0.60 0.60
Table 6: Comparison of F1 scores (%) across five benchmarks. “QR” means the Query Rewriting. “GPT-4o R” means the GPT-4o Recognition.
Results on Medicament Benchmark. Table 3 provides the exploration results of different models on composite medication intents. (i) Attributed to the drug knowledge injection and prototype-guided embedding in MedAide, the reliability of medication counseling responses is consistently improved across all baselines. (ii) In this case, Baichuan4 and GPT-4o demonstrate the most significant overall gains. For example, the GLEU scores separately increase 28.91% and 28.48%. This finding indicates that MedAide facilitates LLMs with large-window contexts to comprehend rich drug information and provide specialized dosage recommendations. (iii) In the zero-shot reasoning pattern, Meditron-7B’s ROUGE-1/2/L remarkably enhances from 12.29/2.12/6.31% to 34.84/29.43/32.57%, verifying that our decision-analysis module provides favorable factual evidence in drug contraindication understanding.
Results on Post-Diagnosis Benchmark. (i) In the post-diagnosis applications from Table 4, we observe that most MedAide-based models are superior to vanilla baselines by large margins. (ii) GPT-4o performs even better in generating personalized rehabilitation recommendations, with a relative improvement of 39.23% in the BLEU-1 score, providing hopefully more reliable rehabilitation guidelines for patients. (iii) Meanwhile, Baichuan4 improves on the ROUGE-L metric by 21.35%, revealing the MedAide-based version’s advantages in processing dynamic medical data and producing progress tracking reports.
4.4 Expert Doctor Evaluation
Expert evaluations play a key role in the practical applications of medical models. We invite 6 doctors (each paid \$300) to select the winners of responses generated by different models before and after the introduction of MedAide by majority voting rule. The response content is holistically evaluated by considering factual accuracy, recommendation practicality, and humanistic care. (i) As shown in Figure 2, all the MedAide-based models exhibit more win rates in different benchmark tests, indicating the effectiveness and applicability of the proposed framework. (ii) MedAide not only improves the healthcare specialization of medical LLMs with different sizes, but also significantly enhances the coping ability of general-purpose models when dealing with complicated medical tasks.
4.5 Ablation Study
We perform systematic ablation studies to explore the effects of different components in Table 5. GPT-4o is chosen as the baseline to combine MedAide.
Necessity of Query Rewriting. Firstly, we remove the query rewriting procedure to assess its impact on performance. (i) The observed significant declines in all metrics indicate that query rewriting is crucial for ensuring that the input information is accurately interpreted by downstream modules with respect to intents. (ii) Additionally, incorporating factual information as context enhances the model’s ability to accurately comprehend user queries, mitigating hallucinations and reducing the generation of ambiguous interpretations.
Importance of Intent Recognition. We replace the learning-based contextual encoder with the prompt-based GPT-4o to explore the impact of different intent recognition strategies. (i) Despite the improvement over the vanilla baseline, GPT-4o recognition is a sub-optimal solution. We argue that the encoder trained with explicitly supervised information can align more purposefully to actual medical intents, producing more personalized and representative judgments than prompt engineering. (ii) Also, our default strategy has better flexibility and extensibility, which can be dynamically optimized according to specific healthcare scenarios.
Effectiveness of Decision Analysis. Furthermore, we directly assemble the different outputs produced by the activated agents to serve as an alternative candidate. The results from the bottom of Table 5 show consistent performance drops of the model on different medical intent understandings, proving the effectiveness of our method. A plausible explanation is that the decision analysis module not only summarizes the different agent outputs in an organized manner, but also provides comprehensive and accurate conclusions based on medical guidelines and patient histories.
Benchmarks MedAide (Ours) MedAgent
Pre-Diagnosis 15.93/17.64 14.33/15.26
Diagnosis 27.80/23.67 25.46/19.54
Medicament 45.13/52.17 28.69/36.07
Post-Diagnosis 26.83/18.12 25.45/16.69
Table 7: Comparison of BLEU-1/ROUGE-L scores (%) between MedAide and MedAgent across benchmarks.
Figure 3: Response visualization analysis on different models for the same user query.
4.6 Intent Detection Analysis
To further observe the intent detection performance, Table 6 compares the F1 scores computed between the intents recognized by MedAide and ground truths under different strategies. We randomly sample 100 instances from each benchmark to compose a more challenging Intent Aggregation benchmark. The core observations are as follows. (i) The complete framework achieves the best results, reaching even 86% performance on the Aggregation benchmark, exhibiting omni intent semantic understanding capabilities. (ii) Significant performance deteriorations occur at “w/o QR” and “w/ GPT-4o R”, implying the effectiveness of fine-grained decomposition of medical queries and the development of our tailored intent recognition mechanism.
Figure 4: The effect of different agent numbers on performance across the four medical benchmarks.
4.7 Collaboration Framework Comparison
Here, we compare the reproducible multi-agent framework MedAgent Tang et al. (2024). This medical framework mainly sets up four department-specific experts, including pediatrics, cardiology, pulmonology, and neonatology. Each expert is based on the GPT-4o implementation and works together to collaboratively generate a unified report as the final result. (i) From Table 7, our framework outperforms MedAgent on all four benchmarks. For instance, in the Medicament benchmark, MedAide beats MedAgent by large margins with absolute gains of 16.44/16.70% on the BLEU-1/ROUGE-L scores. This strength comes from our specialized medication retrieval, which provides a refined context of drug information for the model. (ii) Also, the benefits on Pre-Diagnosis and Post-Diagnosis tasks reflect that the proposed MedAide captures the key characteristics of diagnostic and nursing demands more efficiently.
4.8 Response Visualization Analysis
To intuitively compare the response quality of the different models, we show in Figure 3 the generated results of MedAide and SOTA methods under the same user query. HuatuoGPT-II provides extensive medical knowledge but lacks targeted recommendations. Although the MedAgent framework gives detailed treatment solutions through multidisciplinary perspectives, it falls short of coherent diagnostic reasoning. In contrast, our framework interprets patient outcomes in detail and provides personalized diagnoses and subsequent treatments. MedAide shows favorable usefulness in electrolyte imbalance and cardiovascular risk management.
4.9 Effect of Agent Number
Figure 4 illustrates the variation trend of model performance when the number of agents in the proposed MedAide changes. When the agent number is increased from 2 to 4, there are progressive improvements in the results on all benchmarks, suggesting that multi-agent collaboration helps to improve the model’s performance in multi-tasking scenarios. Conversely, when the agent number exceeds 4 does not bring significant gains due to introducing too much information redundancy.
5 Conclusion and Future Work
This paper introduces MedAide, a multi-agent framework for complex medical scenarios. Leveraging query rewriting, intent recognition, and agent collaboration, MedAide enhances the model’s comprehension of medical intents and demonstrates effectiveness across multiple benchmarks.
Future Work. We plan to incorporate complementary modalities (e.g., medical imaging) in the future to improve the potential of MedAide in multimodal diagnostics and applications.
Limitations
The MedAide framework, while making significant strides in integrating large-scale medical agents with real clinical environments, has its limitations. Currently, the framework incorporates 26,684 drug samples and 506 genuine clinical case records. Despite this scope, it lacks comprehensive coverage, particularly in research on rare diseases and specialized medications. Furthermore, as the initial versions primarily emphasized linguistic processing, future research aims to extend into multimodal capabilities, particularly the integration of medical imaging data, to enhance the framework’s ability to process visual information in clinical diagnostics. The reliance on OpenAI’s API may pose potential operational challenges, suggesting that future studies should explore using more efficient open-source models as viable alternatives. Despite these limitations, the MedAide framework provides a solid foundation for advancing the integration of artificial intelligence in clinical diagnostics.
Ethics Consideration
Ethical considerations are paramount in the application of medical agents to real-world clinical settings. We are fully aware of the potential impacts of our research and have taken deliberate actions to address these issues. To enhance transparency, we are committed to publicly making the drug data and medical records used in our study accessible. This will enable other researchers to validate our findings and build upon our work, fostering collaboration and advancement in this field.
We are acutely aware of the necessity for privacy and data protection. All data utilized has undergone thorough de-identification, with all sensitive information removed, and verified by a partnering medical institution. We invite doctors to perform only evaluations of model responses without involving any form of human subject research. All participants are compensated \$300 for their work, which strictly adheres to the minimum hourly rate for the region in which the work is performed. For the utilization of healthcare-related data, we strictly follow the license agreements of publicly available databases. For the constructed data, we have undergone an internal ethical review by the ethics review board of our partnering medical institutions and are licensed and approved.
Acknowledgment
This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0113502, in part by the Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0103, and in part by the Shanghai Municipal Science and Technology Committee of Shanghai Outstanding Academic Leaders Plan (No. 21XD1430300).
Appendix A Implementation Details
A.1 Query Normalization and Parsing
Algorithm 2 Syntactic Parsing Algorithm
0:  Input query $Q$
0:  Parsed syntactic tree $T$
  Step 1: Tokenize the query
  Let $Q=\{w_{1},w_{2},\dots,w_{n}\}$ be the sequence of words (tokens) in the query.
  Step 2: Initialize parsing
  Initialize an empty syntactic tree $T=\emptyset$.
  Step 3: Apply CFG (Context-Free Grammar)
  for each non-terminal symbol $A$ in grammar $G$ do
     Find a rule $A\rightarrow\alpha$, where $\alpha$ is a string of terminals or non-terminals.
     if $A\rightarrow\alpha$ matches a substring of $Q$ then
        Add this production rule to the tree $T$.
        Replace the matched string with non-terminal $A$.
     end if
  end for
  Step 4: Recursion
  while there are still non-terminals to be parsed do
     Repeat step 3 until $T$ covers the entire query $Q$.
  end while
  Step 5: Output the syntactic tree
  return  $T$
Algorithm 2 first disambiguates the input query and divides the query into individual words or tokens. Next, the algorithm initializes an empty syntax tree and matches substrings in the query by Context-Free Grammar (CFG) rules, adding the conforming rules to the syntax tree. The algorithm keeps applying the rules recursively until the syntax tree covers the entire query and finally outputs the complete syntax tree.
A.2 Refined Query Construction
In the refined query constructor, we implement a set of rules (as shown in Figure 5) to ensure that the generated subqueries are of high quality both semantically and formally. These rules encompass subquery filtering, semantic overlap removal, query consolidation, grammatical normalization, intent prioritization, and format standardization. Together, these processes streamline the output, optimize the query structure, and prioritize critical medical intents, ensuring the final query is both semantically coherent and structurally consistent.
Figure 5: Illustration of the rule sets.
A.3 Training Details of BioBert Encoder
For the medical intent classification, we use the BioBERT-based contextual encoder from the HuggingFace library***https://huggingface.co/dmis-lab/biobert-v1.1. to map the input sentences to a 768-dimensional embedding space. For each intent classification task, we add a fully connected layer after the embedding layer with an output dimension of 17 and generate probability distributions corresponding to the 17 intents via a softmax activation function. To handle these tasks, we deploy four V100 32GB GPUs to provide sufficient computational resources and graphics memory. The Adam optimizer is used in the training process with an initial learning rate of 2e-5, a batch size of 32, and 5 training rounds, and a learning rate scheduling strategy is applied in the validation set for dynamic adjustment. The loss function is the cross-entropy loss function. We evaluate the model by accuracy, precision, recall, and F1 score to ensure the performance of the model on the validation set.
Appendix B Prompt Templates
This section provides an overview of the structured prompt templates employed across four distinct agents within our framework. Each prompt is designed to guide the language model in generating contextually relevant and accurate responses tailored to specific medical tasks. These templates are instrumental in ensuring that the interactions between the agents and users are effective and medically informative. Additionally, the prompt templates for the various components of the query rewriting phase are detailed in Figure 6. The specifics for each agent’s prompt templates are depicted in Figures 7,  8,  9, and 10.
Figure 6: Illustration of the prompt template for the query rewriting.
Figure 7: Illustration of the prompt template for the Pre-Diagnosis.
Figure 8: Illustration of the prompt template for the Diagnosis.
Figure 9: Illustration of the prompt template for the Medicament.
Figure 10: Illustration of the prompt template for the Post-Diagnosis.